OpenStudy (anonymous):

the density of iron is 7.9g/mL and the density of aluminum is 2.7g/mL. what are the volumes of both blocks of iron if each of them has a mass of 5000g?

hero (hero):

v = m/d

OpenStudy (sasogeek):

volume of iron with mass 5000g v=5000/7.9 v=632.911

hero (hero):

(5000g/7.9g)mL (5000g/2.9g)mL

OpenStudy (sasogeek):

volume of aluminum with mass of 5000g v=5000/2.7 v=1851.85mL